a) Two point charges of q1=+3nC and q2=-4nC are separated by a distance of 60 cm. Determine the electric field strength at a point that is 15 cm from q1 and 20 cm from q2

The point lies between the two charges

b) What is the magnitude of a test charge that experiences a force of 1.4 x 10^-8 N at a point where the electric field intensity is 2 x 10^-4 N / C ?


1
Expert's answer
2021-10-03T13:24:24-0400

(a) At 15 cm, the field is

"E_1=\frac{kq_1}{r_1^2}+\frac{kq_2}{(r-r_1)^2}=1376\text{ V/m}."

At 20 cm from the second charge, it is


"E_2=\frac{kq_1}{(r-r_2)^2}+\frac{kq_2}{(r_2)^2}=1067\text{ V/m}."

(b) The charge is


"q=\frac FE=70\space\mu\text{C}."